# Gradient elasticity in Swift-Hohenberg and phase-field crystal models

### v1

@@ 4.2 Crystalline phases
Figure 2: Characteristic lengths $\ell_{1}$ and $\ell_{2}$ by varying $B$ with $C=-1/2$ and $D=1/3$ as from Table 1. Dashed lines correspond to the length scales for one mode approximations, available for triangular and bcc symmetry only.
The values of $\ell_{1,2}$ in Fig. 2 are reported in terms of the ratio with the spatial length scale entering the free energy, namely the lattice parameter $a_{0}=2\pi/k_{0}$, with $k_{0}$ depending on the lattice symmetry (see specific values together with the definition of reciprocal space vectors). They all correspond to fractions of the lattice spacing. Therefore they are in the order of $\sim 10^{-10}$m (Å) when looking at physical units of common materials, in line with GE literature [17, 74]. From a quantitative point of view, however, we note that the ratio of the GE length scales with the lattice parameters does not precisely match typical values computed with ab initio approaches [17]. As with the determination of elastic constants for the classical linear elasticity terms, the considered minimal models encode general features of GE while quantitative descriptions should be addressed by more advanced formulations [57, 75], outlined in the following section.
An important aspect emerging from this analysis is that the GE length scales, being of the order of 1Å, are typically smaller than the coarse-graining lengths that should be considered in PFC-like approaches to work with continuous elastic field (filtering out microscopic fluctuation) or underlying the derivation of the corresponding equations for APFC approaches. Therefore, while GE is self-consistently encoded, specific features emerging at those length scales are expected to be coarse-grained. This aspect is discussed further below with the aid of numerical examples.
@@ 4.3 Characteristic lengths from structural PFC: triangular symmetry
Energy functionals generalizing Eq. (31), or Eq. (33) for amplitude formulations, can be devised by modeling the correlation function in Eq. (29) directly [57, 76]. Such an approach proved powerful in devising parametrization to model specific materials, e.g., in Ref. [75]. We outline here the possibility of devising a parametrization of GE length scales by designing the correlation function. For illustrative purposes, we focus here on the 1-mode triangular symmetry.
For a generic two-point correlation function $C_{2}$, the elastic energy takes the form:
@@
$\mathcal{F}[\psi(\mathbf{r})]=\int\psi(\mathbf{r})\cdot(C_{2}\ast\psi)(\mathbf{r})\,\mathrm{d}^{2}\mathbf{r}.$ (55)
@@
where we consider even terms only due to symmetry consideration, consistently with Eq. (29). For an $N$-term expansion, we then retain up to order $2N$ with $N$ coefficients.
Here we consider a setting that introduces additional parameters while reproducing the features of GE elasticity obtained by the minimal SH energy functional,. Therefore we impose constraints on the form of $\widehat{C_{2}}(|\mathbf{k}|)$ in order to fulfill the following requirements: i) peak in the correlation function at $|\mathbf{k}|=1$ (as above); ii) value of the peak set to 0 such to have the same phenomenological temperature ($B$, see also Sect. 3) as in the minimal models; iii) same elastic constants as the triangular 1-mode (see Table 1); iv) isotropic GE as obtained in Sect. 3.3. All these requirements are independent. Moreover, the last one enforcing isotropic GE imposes two constraints with a total of 5 conditions on coefficients in Eq. (56). Therefore, we need a minimum of $N=6$ to tune the characteristic lengths independently, yielding
$c_{2}=2-3c_{12},\quad c_{4}=1-9c_{12},\quad c_{6}=-8c_{12},\quad c_{8}=0,\quad c_{10}=3c_{12}.$ (57)
This then results in the following length scales, now featuring a free parameter ($c_{12}$)
@@
Note that the value, as well as the ratio between length scales, now depends on the free parameter. Finally, to ensure higher frequencies are suppressed relative to lower ones and real positive characteristic lengths, $c_{12}\geq 0$ and $c_{12}\leq 1/18$ must hold, respectively, thus implying $\ell_{1}/\ell_{2}\in[0,1/\sqrt{3}]$. In this setting, we then reproduced the model conveyed by the minimal SH energy functional, for both linear elasticity and GE, with an extended parameterization. Following similar procedures, we envisage that parametrized SH- or PFC-like models can be devised by considering appropriate expansions of $C_{2}$ and constraints.
@@ 5 Effective GE length scales at dislocations
Coarse-graining should be considered in microscopic models to obtain the macroscopic displacement field $\mathbf{u}$, and thus deformation fields (see also Sect. 2). In PFC models, this coarse-graining length is in the order of the lattice parameter [77, 47], while characteristic GE lengths obtained above are significantly smaller. The interplay with the correlation length of phases $W$, a characteristic length incorporated in the SH energy corresponding to the distance over which the system properties are affected by the phase in a given position [78], is expected. This, for instance, recently emerged for the size of the dislocation core in the PFC model. Although they possess the same topological charge (the Burgers vector), dislocations exhibit a core size that scales with $W$ to some extent [79].
Without approximations, linear, nonlinear, and gradient-elasticity contributions must be considered to fully characterize elasticity in the SH and PFC models. In addition, to accurately encode the dependence of physical constants on model parameters, advanced frameworks must be considered, including, e.g., the variation of amplitude magnitudes [45, 80, 81], assumed constants in the analysis of Sect. 3. These and related aspects prevent us from fully deriving analytic or semi-analytic expressions for GE constants, particularly in realistic settings featuring nonuniform strain variations and nontrivial solutions. However, with the aid of numerical simulations, we show here that the elasticity encoded in the considered models matches very well with known predictions from strain-gradient elasticity by using effective characteristic lengths. Moreover, it allows us to characterize additional effects and unveil their dependence on model parameters.
We look in particular at one of the key evidence concerning GE at the microscale, namely the regularization of the elastic field at dislocations [20], which is also of central interest for crystalline systems. We consider the APFC model for simplicity, encoding a natural coarse-graining length in the equation (the lattice parameter [47]). Accordingly, the elastic field can be directly computed from $\eta_{j}$ rather than upon numerical coarse-graining (see also Eq. 63).
Figure 3: Stress field from APFC modeling of an edge dislocation in a triangular lattice (tri-1) compared to the predictions of GE-1 and GE-2 for an isotropic medium. Characteristic lengths, $(\ell_{\rm I},\ell_{\rm II})$ for GE-1 and $(\varrho_{\rm I},\varrho_{\rm II})$ for GE-2, are obtained by fitting Eqs. (64)–(66) and (67)–(69), respectively, to the APFC numerical result (see further details and values obtained by fitting in C). APFC results are shown with solid grey contour lines and the underlying colormap to note negative and positive stress regions. GE-1 with fitted length scales is illustrated via dot-dashed blue lines, and GE-2 via dashed red lines. Contour lines are set to the same values in all the 2D plots of the stress to $[-7,-3.5,0.0,3.5,7]\times 10^{-3}$. The same applies to the range of color maps. (a) 2D contour plots of the stress field components $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{xy}$ (from left to right) for $A=1$, $B=0.02$. (b) Detailed comparison as 1D curves of the values in panel (a) along $y$ at $x=0$ for $\sigma_{xx}$ and $\sigma_{yy}$, and along $x$ for $\sigma_{xy}$ (from left to right). (c) $\sigma_{xx}$ illustrated as in panel (a) for three different parameter choices. From left to right parameters $(A,B)$ are set to: $(0.5,0.02)$, $(2.0,0.02)$ and $(1,-0.026)$. In all plots $b_{x}=-a_{0}$, $C=-1/2$, and $D=1/3$.
@@
the dislocation displacement field, $\mathbf{b}=b_{x}\hat{\mathbf{x}}$ the Burgers vector, and $\nu$ the Poisson’s ratio [24]. In general, defect arrangements evolve dynamically [44]. To study elastic fields at defects, we consider a static configuration. It consists of a checkered pattern, nominally on a square grid, of dislocations having ±$\pm\mathbf{b}$ as Burgers vectors and distance $L/2$. Notice, however, that due to the underlying triangular symmetry, dislocations cannot be placed exactly on a square grid, and the relaxation of the initial condition leads to a small shift in their positions. However, this configuration allows simulating a static, periodic system by considering an $L\times L$ box. We remark that the displacement field in Eqs. (59) is exact for an isotropic medium within linear elasticity [24]. Here, it is exploited to introduce dislocation with the desired Burgers vector. By letting the system relax to equilibrium, its adaptation to the actual solution, including the effects outlined in (37), is obtained. Analogous results, but with a slower convergence to the equilibrium solution, can be obtained by setting proper singularities in $\theta_{n}$ without initial assumption on the elastic field as done, e.g., in Refs. [48]. Numerical results are obtained by exploiting a simple pseudo-spectral Fourier method as summarized in A.
@@
Figure 3 illustrates the stress field of an edge dislocation in a triangular crystal as obtained at equilibrium by APFC simulations compared to the prediction of GE theories with fitted characteristic lengths. We consider solutions for an edge dislocation in an isotropic medium obtained within GE-1 and GE-2 as reported in Refs. [21] and [23], respectively. Corresponding equations are reported in B, while details of the fitting procedure are given in C (fitted values of the characteristic lengths $\ell_{i}$ and $\varrho_{i}$ are indicated with the subscripts “I” and “II” in the following). GE-1 is the theory emerging from the analysis reported in Sect. 3. GE-2 includes higher-order contributions. Importantly, the solution of the elastic field for a dislocation in GE-2 features smaller gradients (see Fig. 5) at the core, as further discussed below.
@@
In particular, Fig. 3a shows stress field components in 2D for a selected set of parameters. Color maps and representative grey contour lines illustrate the value obtained by numerical APFC simulations. The stress field along lines crossing the dislocation core parallel to the $y$-axis for $\sigma_{xx}$ and $\sigma_{yy}$ and the $x$-axis for $\sigma_{xy}$ are shown in Fig. 3b. $\sigma_{ij}$ from APFC simulations are regularized at the core, and both GE-1 and GE-2 solutions match very well the results of numerical simulations for some values of the respective characteristic lengths ($\ell_{i}$ and $\varrho_{i}$). As observed in other works [47], the elastic field obtained by APFC is asymmetric, a feature that is absent in the considered GE theories as it results from nonlinearities in the free-energy functional [45]. Accordingly, this effect is more pronounced for larger stress values, namely for $\sigma_{xx}$ where the stress field reaches maximum and minimum values two times larger than other components’. Length scales have been determined by fitting the numerically computed values for the stress components via the analytical solution for both GE-1 and GE-2 (reported in B). Characteristic lengths and dislocation positions (subject to the small shift mentioned above) are retained as fitting parameters. Their values, determined via the fitting performed to obtain the plots in Fig. 3a and Fig. 3b, are fully reported in Table 2.
@@
From a quantitative point of view (see Table 2), characteristic GE-1 lengths obtained from fitting $\sigma_{xx}$ and $\sigma_{yy}$ are very similar, $\ell_{\rm I}\sim\ell_{\rm II}\sim 0.55a_{0}$. The values for $\sigma_{xy}$ deviate from this behavior ($\ell_{\rm I}\sim 0.23a_{0}$ and $\ell_{\rm II}\sim 0.48a_{0}$). Nevertheless, by using the value obtained for the other components, $\sigma_{xy}$ actually shows a very similar profile with respect to the maximum and minimum values as well as the decay far from the core. A small difference is observed only very close to the core, where the two significantly different lengths allow for smaller gradients and a better matching of the simulation results. When enforcing a single characteristic length ($\ell_{1}=\ell_{2}$)—which we note in passing would again be consistent with the conclusion of lattice theory with central force interaction between atoms [72, 67]—the fitted values are very similar (see the third line in Table 2). Therefore, this deviation results from the larger number of degrees of freedom in the model and shows that GE-1 can match the simulations well but with characteristic lengths that are not fully consistent among different stress-field components. Conversely, the values obtained via fitting the characteristic GE-2 lengths are very close to each other in all three stress components, assessing that the underlying model accurately describes simulation results. By relaxing the assumption of constant magnitude of the amplitudes $\eta_{j}$ underlying the derivation in Sect 3.3, we indeed find that additional contributions appear, including terms entering GE-2. Further details to substantiate this argument are reported in D.
@@
Figure 4: Variation of characteristic lengths with (a) $A$ for $B=0.02$, and (b) $B$ for $A=1$. In all simulations, $C=-1/2$ and $D=1/3$. The panels illustrate in particular log-log plots of ¯$\bar{\ell}=(\ell_{\rm I}+\ell_{\rm II})/2$ (blue circles), $\omega$ (red squares) and $\gamma$ (orange triangles). Solid lines are guidelines for the eyes. Values reported here are obtained by fitting $\sigma_{yy}$. Dependence on $B$ is illustrated via the quenching depth $|B-B_{c}|$. The correlation length $W$ from Eq. (70) is reported for comparison (dashed line).
@@
The agreement between the simulated and analytical GE stress fields, as well as more detailed considerations above, is found to be valid across a broad parameter range. This is first shown in Fig. 3c where we report 2D stress maps and contour lines of $\sigma_{xx}$ for three simulations featuring $A$ or $B$ values different from those in Fig. 3a. The change in the $\sigma_{xx}$ is reflected in a variation of the characteristic lengths determined by fitting. This aspect is further quantitatively illustrated in Fig. 4, also allowing for further comparisons between the considered GE theories. Therein, we discuss the values of fitted characteristic lengths from $\sigma_{yy}$. This component is chosen as it features an almost symmetric profile owing to lower maximum values for the stress field and thus a limited contribution of nonlinearities compared to $\sigma_{xx}$ (see also Fig. 3). For GE-1 we show the average of the two characteristic lengths, ¯$\bar{\ell}=(\ell_{\rm I}+\ell_{\rm II})/2$. Consistent with the values in Table 1 (for the simulation illustrated in Fig. 3a-b), the relative difference between their values is however below $5\%$ over the whole range of parameters considered. For GE-2 we report the two characteristic lengths $\omega=\sqrt{\varrho_{\rm I}^{2}+\varrho_{\rm II}^{2}}$ and $\gamma=\sqrt[4]{\varrho_{\rm I}^{2}\varrho_{\rm II}^{2}}$, allowing for direct comparisons with GE-1. We vary in particular parameters $A$ (Fig. 4a) and $B$ (Fig. 4b). The variation with $B$ is illustrated in terms of the quenching depth, namely $|B-B_{c}|$ with $B_{c}=8C^{2}/(135D)$ the value of $B$ at which ordered (periodic $\psi$ with triangular symmetry) and disordered (constant $\psi$) phases have the same energy [47]. We note that the parameter range includes most variations typically considered for PFC models.
@@
We generally find that effective GE lengths are larger than the one computed in Sect. 3 and also better approach common values reported in the literature for GE length scales [17, 74]. They approach values obtained in Sect. 3 for small $A$ and $B$. Characteristic lengths of GE-1 and GE-2 vary similarly. Interestingly, $\omega\sim 0.9\bar{\ell}$. We remark that $\omega$ is the characteristic length to the lowest order of GE-2, which reduces to GE-1. Still, these values do not match exactly and, indeed, $\gamma\neq 0$. Therefore, we assess further the presence of higher-order corrections from terms peculiar to GE-2, as suggested by the argument in D.
@@
Importantly, the characteristic lengths are found to be proportional to the correlation length $W$. In the considered models, this quantity can be determined as the width of the interface between an ordered and disordered phase, both in equilibrium and nonequilibrium settings [78]. For the one-mode approximation of the triangular lattice, it has an analytic solution, reported in Eq. (70). The same or slightly varying proportionality constant over the whole range of explored parameters is obtained, namely ¯$\bar{\ell}/W\sim{\omega}/W\in(0.4,0.6)$. We remark, however, that the expression of $W$ used here, Eq. (70), is also derived in an approximated setting. In models like SH or PFC, the extension of the defect core is also found to scale similarly with the correlation length [79]. Therefore, the effective characteristic length is found to scale de facto with the core size, consistent with continuum descriptions of the elastic field of dislocations in GE theories [20, 21, 23]. Importantly, by recalling that $B$ is a phenomenological temperature parameter, the effective GE characteristic lengths at dislocations are then found to be connected to the temperature.
6 Conclusion

### v2

@@ 4.4 Discussion of the results
Figure 2: Characteristic lengths $\ell_{1}$ and $\ell_{2}$ by varying $B$ with $C=-1/2$ and $D=1/3$ as from Table 1. Dashed lines correspond to the length scales for one-mode approximations, which are only available for triangular and bcc symmetry.
The values of $\ell_{1,2}$ in Fig. 2 are reported in terms of the ratio with the spatial length scale entering the free energy, namely the lattice parameter $a_{0}=2\pi/k_{0}$, with $k_{0}$ depending on the lattice symmetry (see specific values together with the definition of reciprocal space vectors in Sect. 4.3). They all correspond to fractions of the lattice spacing. Therefore, they are in the order of $\sim 10^{-10}$m (Å) when looking at physical units of common materials, in line with GE literature [19, 78]. From a quantitative point of view, as we may expect considering the coarse-grained nature of SH and PFC models, we note that the ratio of the GE length scales with the lattice parameters does not precisely match typical values computed with ab initio approaches [19]. Suitable extensions by considering more advanced formulations [79, 80] are outlined in the following section.
An important aspect emerging from this analysis is that the GE length scales ($\sim$ 1Å) are typically smaller than the coarse-graining length that should be considered in PFC-like approaches to work with continuous elastic fields (filtering out microscopic fluctuation) or that underlies the derivation of the equations for APFC approaches. Therefore, while GE is self-consistently encoded, specific features emerging at those length scales are expected to be coarse-grained. This aspect is discussed further below with the aid of numerical examples.
@@ 4.5 Characteristic lengths from structural PFC: triangular symmetry
Energy functionals generalizing Eq. (28), or Eq. (30) for amplitude formulations, can be devised by considering correlation function instead of the differential operator [79, 81]. Such an approach proved powerful in devising parametrizations to model specific materials, e.g., in Ref. [80]. Here, we outline the possibility of parametrizing the GE length scales by designing the correlation function. For a generic two-point correlation function $C_{2}$, the elastic energy takes the form:
@@
$\mathcal{F}[\psi(\mathbf{r})]=\int\mathrm{d}\mathbf{r}\,\bigg{\{}\psi(\mathbf{r})\int\mathrm{d}\mathbf{r}^{\prime}\bigg{[}C_{2}(\mathbf{r}-\mathbf{r}^{\prime})\psi(\mathbf{r}^{\prime})\bigg{]}\bigg{\}}.$ (53)
@@
where we only consider even terms due to symmetry considerations. For an $N$-term expansion, the highest order is thus $2N$.
We consider a setting that introduces additional parameters while reproducing the features of GE obtained by the minimal SH energy functional, focusing on the 1-mode triangular symmetry for illustration purposes. We thus impose constraints on the form of $\widehat{C_{2}}(|\mathbf{k}|)$ to fulfill the following requirements: (i) peak in the correlation function at $|\mathbf{k}|=1$ (as above); (ii) value of the peak set to 0 such to have the same phenomenological temperature/quenching depth ($B$, see also Sect. 3) as in the minimal models; (iii) same elastic constants as the triangular 1-mode (see Table 1); (iv) isotropic GE as obtained in Sect. 4.1. As these requirements are independent, and enforcing isotropic GE imposes two additional constraints, the coefficients in Eq. (54) must satisfy a total of 5 conditions. Therefore, a minimum of $N=6$ is needed to tune the characteristic lengths independently, yielding
$c_{2}=2-3c_{12},\quad c_{4}=1-9c_{12},\quad c_{6}=-8c_{12},\quad c_{8}=0,\quad c_{10}=3c_{12},$ (55)
which results in the following length scales, now featuring a free parameter ($c_{12}$)
@@
Note that the value and the ratio between length scales now depend on $c_{12}$. Finally, to ensure the suppression of higher frequencies and real positive characteristic lengths, $c_{12}\geq 0$ and $c_{12}\leq 1/18$ must hold, respectively, thus implying $\ell_{1}/\ell_{2}\in[0,1/\sqrt{3}]$. In this setting, we then obtained a SH energy functional with an extended parameterization, allowing for tuning GE constants. Following similar procedures, we envisage that extended parametrization can be devised for models based on a SH free energy encoding other lattice symmetries or coupling with other terms by considering appropriate expansions of $C_{2}$.
@@ 5 Effective GE length scales at dislocations
To fully characterize elasticity in the SH and PFC models, linear, nonlinear, and strain-gradient contributions must be considered. Consequently, the assumptions considered in Sect. 4 to allow for the derivation of fully analytic expressions need to be relaxed. Through numerical simulations, we show in this section that the elasticity encoded in the considered models matches well with known predictions from strain-gradient elasticity using effective characteristic lengths. These are found to be larger than the analytical predictions reported above and thus are in better agreement with results from the literature. Moreover, it allows us to characterize additional effects and unveil their dependence on model parameters.
5.1 Numerical simulation of a stationary defect configuration
A key evidence concerning GE at the microscale is the regularization of the elastic fields at dislocations [22], which is also of central interest for crystalline systems. For simplicity, we consider the APFC model from Sect. 3.2, thus encoding a natural coarse-graining length in the equation (the lattice parameter [49]). Accordingly, the elastic field can be directly computed from $\eta_{j}$ rather than upon numerical coarse-graining (see also Eq. 61).
@@ 5.1 Numerical simulation of a stationary defect configuration
the dislocation displacement field, $\mathbf{b}=b_{x}\hat{\mathbf{x}}$ the Burgers vector, and $\nu$ the Poisson’s ratio [27]. In general, defect arrangements evolve dynamically [46]. To study elastic fields at defects, we consider a static configuration. It consists of a checkered pattern, nominally on a square grid, of dislocations having ±$\pm\mathbf{b}$ as Burgers vectors and distance $L/2$. Notice, however, that dislocations cannot always be placed exactly on a square grid, e.g. in the presence of a triangular lattice, and the relaxation of the initial condition may lead to a small shift in their positions. However, this configuration allows simulating a static, periodic system by considering an $L\times L$ box. We remark that the displacement field in Eqs. (57) is exact for an isotropic medium within linear elasticity [27]. Here, it is exploited to introduce dislocation with the desired Burgers vector. By letting the system relax to equilibrium, the actual solution including the effects outlined in Eq. (34), is obtained. Analogous results, but with a slower convergence to the equilibrium solution, can be obtained by setting proper singularities in $\theta_{n}$ without initial assumption on elasticity as done, e.g., in Ref. [50]. Numerical results are obtained by exploiting a simple pseudo-spectral Fourier method as summarized in A.
5.2 Elastic field at a dislocation
Figure 3: Stress field from APFC modeling of an edge dislocation in a triangular lattice (tri-1) compared to the predictions of GE-1 and GE-2 for an isotropic medium. Characteristic lengths, $(\ell_{\rm I},\ell_{\rm II})$ for GE-1 and $(\varrho_{\rm I},\varrho_{\rm II})$ for GE-2, are obtained by fitting Eqs. (62)–(64) and (65)–(67), respectively, to the APFC numerical result (see further details and values obtained by fitting in C). APFC results are shown with solid grey contour lines and the underlying colormap to note negative and positive stress regions. The GE-1 solution with fitted length scales is illustrated via dot-dashed blue lines, and the GE-2 solution via dashed red lines. Contour lines are set to the same values in all the 2D plots of the stress to $[-7,-3.5,0.0,3.5,7]\times 10^{-3}$. The same applies to the range of color maps. (a) 2D contour plots of the stress field components $\sigma_{xx}$, $\sigma_{yy}$, and $\sigma_{xy}$ (from left to right) for $A=1$, $B=0.02$. (b) Detailed comparison as 1D curves of the values in panel (a) along $y$ at $x=0$ for $\sigma_{xx}$ and $\sigma_{yy}$, and along $x$ at $y=0$ for $\sigma_{xy}$ (from left to right). (c) $\sigma_{xx}$ illustrated as in panel (a) for three different parameter choices. From left to right parameters $(A,B)$ are set to: $(0.5,0.02)$, $(2.0,0.02)$ and $(1,-0.026)$. In all plots $b_{x}=a_{0}$, $\bar{\psi}=0$, $C=-1/2$, and $D=1/3$.
@@ 5.2 Elastic field at a dislocation
Figure 3 illustrates the stress field of an edge dislocation in a triangular crystal obtained at equilibrium by APFC simulations compared to the prediction of GE theories with fitted characteristic lengths. We consider solutions for an edge dislocation in an isotropic medium obtained within GE-1 and GE-2 as reported in Refs. [23] and [25], respectively. Corresponding equations are reported in B. Details of the fitting procedure are given in C. Fitted values of the characteristic lengths $\ell_{i}$ and $\varrho_{i}$ are indicated with the subscripts “I” and “II” in the following.
@@
In particular, Fig. 3a shows stress field components in 2D for a selected set of parameters. Color maps and representative grey contour lines illustrate the value obtained by numerical APFC simulations. The stress field along lines crossing the dislocation core parallel to the $y$-axis for $\sigma_{xx}$ and $\sigma_{yy}$ and the $x$-axis for $\sigma_{xy}$ are shown in Fig. 3b. $\sigma_{ij}$ from APFC simulations are regularized at the core, and both GE-1 and GE-2 solutions match very well the results of numerical simulations for some values of the respective characteristic lengths ($\ell_{i}$ and $\varrho_{i}$). As observed in other works [49], the elastic field obtained by APFC is asymmetric, a feature that is absent in the considered GE theories as it results from nonlinearities in the free-energy functional [47]. Accordingly, this effect is more pronounced for larger stress values, namely for $\sigma_{xx}$ where the stress field reaches maximum and minimum values two times larger than other components. Characteristic lengths have been determined by fitting the numerically computed values for the stress components via the analytical solution for both GE-1 and GE-2 (reported in B). Dislocation positions (subject to the small shift mentioned above) are retained as fitting parameters too. All the values determined via fitting for the plots in Fig. 3a and Fig. 3b are reported in Table 2.
While our focus here is on the prototypical case of an edge dislocation in an isotropic lattice, which allows for extended comparisons with analytical expressions and, in principle, connections to other theories, it is important to note that other dislocations (like of screw or mixed type) and lattice symmetries are also naturally described in PFC models [39, 82, 83, 74, 46]. An extended study covering more cases is out of the scope of the present work. The analysis of targeted cases is left to future applications.
5.3 Discussion of the results
@@
From a quantitative point of view (see Table 2), characteristic GE-1 lengths obtained from fitting $\sigma_{xx}$ and $\sigma_{yy}$ are very similar, $\ell_{\rm I}\sim\ell_{\rm II}\sim 0.55a_{0}$. The values for $\sigma_{xy}$ deviate from this behavior ($\ell_{\rm I}\sim 0.23a_{0}$ and $\ell_{\rm II}\sim 0.48a_{0}$). Nevertheless, by using the value obtained for the other components, $\sigma_{xy}$ shows a very similar profile with respect to the maximum and minimum values as well as the decay far from the core. A small difference is observed only very close to the core, where the two significantly different lengths allow for smaller gradients and a better matching of the simulation results. When enforcing a single characteristic length ($\ell_{1}=\ell_{2}$)—which would be consistent with the conclusion of lattice theory with central force interaction between atoms [76, 7]—the fitted values are very similar for the three components of the stress field analyzed (see the third line in Table 2). We may conclude that GE-1 can match the simulations well but with characteristic lengths that are not fully consistent among different stress-field components. Conversely, the values obtained via fitting the characteristic GE-2 lengths are very close to each other in all three stress components, assessing that the underlying model accurately describes simulation results. By relaxing the assumption of constant amplitude moduli underlying the derivation in Sect 4.1, we indeed find that additional contributions appear, including terms entering GE-2. Further details to substantiate this argument are reported in D.
@@
Figure 4: Dependence of the characteristic lengths on free energy parameters. Variation of characteristic lengths with (a) $A$ for $B=0.02$, and (b) $B$ for $A=1$. In all simulations, $C=-1/2$ and $D=1/3$. The panels illustrate in particular log-log plots of ¯$\bar{\ell}=(\ell_{\rm I}+\ell_{\rm II})/2$ (blue circles), $\omega$ (red squares) and $\gamma$ (orange triangles). Solid lines are guidelines for the eyes. Values reported here are obtained by fitting $\sigma_{yy}$. Dependence on $B$ is illustrated via the quenching depth $|B-B_{c}|$. The correlation length $W$ from Eq. (68) is reported for comparison (dashed line).
@@
The agreement between the simulated and analytical stress fields holds for a broad parameter range. This is first shown in Fig. 3c, via 2D stress maps and contour lines of $\sigma_{xx}$ for three simulations featuring $A$ or $B$ values different from those in Fig. 3a. The change in the $\sigma_{xx}$ is reflected in a variation of the characteristic lengths determined by fitting. This aspect is further quantitatively illustrated in Fig. 4, allowing for a comprehensive analysis of the considered GE theories. For this analysis we consider the characteristic lengths fitted from $\sigma_{yy}$. This component is chosen as it features an almost symmetric profile owing to lower maximum values for the stress field, and thus a limited contribution of nonlinearities, than $\sigma_{xx}$ (see also Fig. 3). For GE-1, we show the average of the two characteristic lengths ¯$\bar{\ell}=(\ell_{\rm I}+\ell_{\rm II})/2$, as their relative difference is below $5\%$ over the whole range of parameters considered (see Table 1). For GE-2 we report the two characteristic lengths $\omega=\sqrt{\varrho_{\rm I}^{2}+\varrho_{\rm II}^{2}}$ and $\gamma=\sqrt[4]{\varrho_{\rm I}^{2}\varrho_{\rm II}^{2}}$, allowing for direct comparisons with GE-1 (see also Sect. 2.3). We vary in particular parameters $A$ (Fig. 4a) and $B$ (Fig. 4b). The variation with $B$ is illustrated in terms of the quenching depth, namely $|B-B_{c}|$ with $B_{c}=8C^{2}/(135D)$ the value of $B$ at which ordered (periodic $\psi$ with triangular symmetry) and disordered (constant $\psi$) phases have the same energy [49]. We note that the parameter range includes most variations typically considered for PFC models.
@@
We generally find that effective GE characteristic lengths are larger than the one computed in Sect. 4 and approach common values reported in the literature in terms of ratio with the lattice parameter [19, 78]. Characteristic lengths of GE-1 and GE-2 vary similarly. Interestingly, $\omega\sim 0.9\bar{\ell}$. We remark that $\omega$ is the characteristic length to the lowest order of GE-2, which reduces to GE-1 in the limit $\gamma=0$. The computed value of $\gamma$ being greater than zero is thus a further demonstration of the presence of higher-order corrections from terms specific to GE-2, as supported by the argument in D.
@@
Importantly, the characteristic lengths are found to be proportional to the correlation length $W$, which corresponds to the width of the interface between an ordered and disordered phase, both in equilibrium and nonequilibrium settings [84]. Dashed black lines in Fig. 4 show the analytic expression of $W$ reported in Eq. (68), which is derived in an approximated setting similar to the ones considered in Sect. 4. Ultimately, we find a proportionality constant ¯$\bar{\ell}/W\sim{\omega}/W\in(0.4,0.6)$. In models like SH or PFC, the extension of the defect core scales similarly with the correlation length [85]. Therefore, the effective characteristic length—which we recall encodes microscopic effects into (continuum) elasticity—is found to scale de facto with the core size, consistently with continuum descriptions of the elastic field of dislocations in GE theories [22, 23, 25]. By recalling that $B$ is a phenomenological temperature parameter, the effective GE characteristic lengths at dislocations are then connected to the temperature.
6 Conclusions